OpenStudy (anonymous):

What is 125 written as a power of a number? A. 5*25 B. 5*2 C. 5*3 D. 2*55

OpenStudy (anonymous):

C

OpenStudy (anonymous):

sure?

OpenStudy (anonymous):

C. 5^3 because when you raise it to a power it means you multiply that number by itself that number of times. So, for 5^3 = 5*5*5 which equals 125

OpenStudy (anonymous):

yeah sure bcoz 5^3 =5*5*5=25*5=125